### v3

# Rates of Convergence in the Central Limit Theorem for Markov Chains, with an Application to TD Learning

Abstract
We prove a non-asymptotic central limit theorem for vector-valued martingale differences using Stein’s method, and use Poisson’s equation to extend the result to functions of Markov Chains. We then show that these results can be applied to establish a non-asymptotic central limit theorem for Temporal Difference (TD) learning with averaging.
1 Introduction
Starting with the seminal work of (Polyak and Juditsky 1992), where they analyzed the averaging rule of (Polyak 1990, Ruppert 1988), the central limit theorem for martingales has been a valuable tool to understand asymptotic efficiency of stochastic approximation algorithms for optimization. More recently, the central limit theorem for functions of Markov chains has also been used to study the asymptotic efficiency of machine learning algorithms (Borkar et al. 2021, Hu et al. 2024). However, in machine learning applications, finite-time (rather than asymptotic) bounds are preferred to understand the sample complexity of algorithms. However, to the best of our knowledge, there is limited work on the central limit theorems for vector-valued martingales and vector-valued functions of Markov chains that would allow one to both quantify asymptotic efficiency and obtain finite-time bounds for learning and optimization algorithms. A recent paper by (Anastasiou et al. 2019) derives a rate of convergence for the central limit theorem vector-valued martingales; however, the notion of distance between probability distributions in that paper is not sufficient for some applications.
Our goal in this paper is to obtain bounds on the rate of convergence to normality to complement the central limit theorem for Markov chains (Meyn and Tweedie 2012). To this end, we first establish a result for martingale differences by combining the Lindeberg decomposition along with Stein’s method, which was first used in (Röllin 2018) for scalar valued martingale differences. This key idea in (Röllin 2018) was extended to the case of vector-valued martingales in (Anastasiou et al. 2019) who also used it to obtain a non-asymptotic version of the central limit theorem in (Polyak and Juditsky 1992). Our approach builds upon the ideas in (Röllin 2018, Anastasiou et al. 2019), but we also introduce new techniques in order to obtain stronger results:
  • We would like our results to hold when the deviation from normality is measured using the Wasserstein distance. The results in (Anastasiou et al. 2019) provide bounds only for weaker distances. Therefore, we have to obtain results for more general test functions than the ones considered in (Anastasiou et al. 2019). To do so, we leverage the results on the regularity of solutions to Stein’s equation for vector-valued random variables in (Gallouët et al. 2018, Fang et al. 2019), which makes our analysis considerably different.
  • To obtain rates of convergence for the Markov chain central limit theorem, we convert a Markov chain into a martingale difference noise using Poisson’s equation; see (Metivier and Priouret 1984, Benveniste et al. 2012, Douc et al. 2018), for example. To the best of our knowledge, this idea has not been used to derive convergence rates for the central limit theorem for functions of Markov chains.
A difference between our results and the results in (Röllin 2018, Anastasiou et al. 2019) for martingales is the fact that we state our convergence bounds in terms of a positive definite matrix $\Sigma_{\infty}$ which can be viewed as the limit of conditional covariances of martingale differences (see Theorem 1 and the short discussion after the theorem). This is important to us since our goal is to state our results for Markov chains in terms of the asymptotic covariance, which is important in applications; see (Borkar et al. 2021), for example. Therefore, our assumptions are stated differently than in (Röllin 2018, Theorem 1) or in (Anastasiou et al. 2019, Equation 3.1). It is also worth remarking that Stein’s method to derive rates of convergence for the martingale central limit theorem is relatively new. The early papers on obtaining rates of convergence for the martingale central limit theorem use other techniques; see, for example, (Bolthausen 1982) and other references in (Röllin 2018). Here, we use Stein’s method because we found it easier to quantify the impact of the asymptotic variance on the rate of convergence using Stein’s method. Specifically, we show that the regularity results in (Gallouët et al. 2018, Fang et al. 2019) and the calculations in (Röllin 2018, Gallouët et al. 2018) can be used, along with insights from Markov chain theory, to understand the impact of the asymptotic covariance term on the rates of convergence in the martingale and Markov chain central limit theorems.
As an application, we apply the Markov chain central limit theorem to study the distribution of error in TD learning. TD learning is the most widely studied algorithm for evaluating the performance of a given policy from data in Markov Decision Processes (Sutton and Barto 2018). The convergence of TD learning with decaying stepsizes was proved in (Tsitsiklis and Van Roy 1996). Over the last few years, there has been a resurgent interest in understanding the non-asymptotic convergence behavior of TD learning and further using it to study other reinforcement algorithms which use TD learning within their framework. The first result of this type was obtained in (Bhandari et al. 2018) who used a projection step in their version of the TD learning algorithm. The first finite-time bounds for unprojected TD learning were obtained in (Srikant and Ying 2019), who studied both fixed step-sizes and decaying step-sizes. TD learning can be viewed as a special case of linear stochastic approximation with Markovian and multiplicative noise and as a special case of Markovian jump linear systems (Hu and Syed 2019). The results in (Srikant and Ying 2019) were extended to nonlinear, contractive stochastic approximation algorithms in (Chen et al. 2023a), who also derived finite-time performance bounds for specific choices of decaying step-size rules, which also allowed them to study the robustness of different step-size choices. Finite-time performance bounds for variants of TD learning and other schemes have been studied in (Gupta et al. 2019, Doan 2021, 2022) and concentration results for TD learning have been obtained in (Telgarsky 2022, Chen et al. 2023b).
While most of the above papers study standard TD learning, in practice, it is well-known that averaging the iterates can help improve the asymptotic variance. This idea was originally studied for stochastic approximation and stochastic gradient descent in the seminal works in (Ruppert 1988, Polyak 1990, Polyak and Juditsky 1992); see (Moulines and Bach 2011) for a finite-time analysis. More recently, the asymptotic behavior of TD learning and other algorithms with averaging has been studied in (Borkar et al. 2021, Mou et al. 2021, Durmus et al. 2022); see also (Li et al. 2023a) for related work. As in the papers on stochastic approximation, the conclusion in (Borkar et al. 2021) is that, for a suitable choice of decaying stepsizes, averaging the output leads to efficient estimators in the sense that the variance is minimized. Further, (Borkar et al. 2021) also established functional central limit theorems and central limit theorems for TD learning and other reinforcement learning algorithms. The work in (Mou et al. 2021) shows order-optimal rate of convergence for the mean-squared error, while (Durmus et al. 2022) obtains higher moment bounds and high probability deviation bounds. In contemporaneous work to this paper, finite-time performance bounds on the covariance of the averaged TD learning iterates have been obtained in (Haque et al. 2023) using more robust step-size rules. Their model is more general; they study two-time scale linear stochastic approximation algorithms, with TD learning with averaging being a special case of the models that they study. A central limit theorem for two time-scale stochastic approximation has also been established for two time-scale stochastic approximation in (Hu et al. 2024). Related work also includes a central limit theorem for Polyak-Ruppert averaged Q-learning (Li et al. 2023b).
Our main results in the paper are as follows:
  • We derive non-asymptotic central limit theorems for martingales and Markov chains by exploiting ideas from (Röllin 2018, Gallouët et al. 2018, Fang et al. 2019).
  • We then extend the results to Markov chains using Poisson’s equation to relate Markov chains to martingales. In this case, the asymptotic covariance can be characterized as is well known in the literature; see (Douc et al. 2018), for example.
  • Finally, we apply the Markov chain results, along with the ideas in (Polyak and Juditsky 1992), to estimate the rate of convergence in the central limit theorem for TD learning with Polyak-Ruppert averaging. An advantage of the approach here is that our result directly establishes a central limit theorem for the scaled averaged of the iterates, instead of first establishing a function central limit theorem for the entire trajectory of the iterates and then establishing the central limit theorem.
  • In addition to the difference in the notion of distance considered here and in (Anastasiou et al. 2019), in the TD learning application, we have to deal with the fact that the noise multiplies the TD learning parameters, whereas the noise is additive in the stochastic gradient application considered in (Anastasiou et al. 2019).
It is important to note that, unlike (Polyak and Juditsky 1992), our result only holds for averaging along with an appropriate decaying step-size rule. Specifically, the result does not hold for fixed step-size TD learning with averaging due to the bias issue pointed out in (Borkar et al. 2021, Lauand and Meyn 2023, Huo et al. 2023)); also see (Nagaraj et al. 2020, Roy and Balasubramanian 2023, Jain et al. 2018) for challenges in dealing with Markov noise in other problems. A nonasymptotic central limit theorem has been established for constant step-size stochastic gradient descent in (Dieuleveut et al. 2020), but the model considered here is different from the problem in (Dieuleveut et al. 2020) (for example, Markovian noise, constant step-size vs decaying step-size) and therefore, the techniques there do not seem to apply to our problem.
The rest of the paper is organized as follows. In Section 2, we establish an upper bound on the rate of convergence in the martingale central limit theorem under some assumptions, and in Section 3, we apply these results to bound the rate of convergence in the central limit theorem for functions of Markov chains using Poisson’s equation. The Markov chain results are then applied to study the central limit for averaged iterates of TD leaning in Section 4. Concluding remarks and suggestions for future work are provided in Section 5.
2 Martingale Central Limit Theorem
In this section, we present a bound on the rate of convergence for the martingale central limit theorem in terms of the Wasserstein distance using Stein’s method.
2.1 Preliminaries
In this subsection, we will describe the notation we use to establish the rate of convergence in the central limit theorem for vector-valued martingale differences. Let $\{m_{k}\}_{k\geq 1}$ be a $d$-dimensional martingale difference sequence with respect to a filtration $\{\mathcal{F}_{k}\}_{k\geq 0}$ i.e., $m_{k}$ is $\mathcal{F}_{k}$-adapted, $E(||m_{k}||)<\infty,$ where |||||$\cdot$||| | ⋅ | | is the standard Euclidean norm in $\Re^{d}$ and
$E(m_{k}|\mathcal{F}_{k-1})=0\qquad\forall k\geq 1.$
Assumption 1.
We assume that the martingale difference sequence satisfies the following properties:
  1. 1.
    $E(||m_{k}||^{2+\beta})$ exists a.s. for all $k\geq 1$ and some $\beta\in(0,1).$
  2. 2.
    $\Sigma_{k}=E(m_{k}m_{k}^{T}|\mathcal{F}_{k-1})$ exists for all $k$.
$\diamond$
Item (1) in the assumption is motivated by the fact that we want to extend the techniques in (Gallouët et al. 2018) to martingales .
Define $S_{n}=\sum_{k=1}^{n}m_{k}.$ We are interested in whether the following normalized sum
$W_{n}:=\frac{S_{n}}{\sqrt{n}}$ (1)
converges to $N(0,\Sigma_{\infty})$, and if so, we are interested in the rate of convergence. We will use Stein’s method to establish such a convergence and to estimate the rate of convergence (Stein 1972, Ross 2011, Chatterjee 2014).
The Wasserstein distance between two $d$-dimensional random vectors $X$ and $Y$ is defined as
$d_{\mathcal{W}}(X,Y)=\sup_{h\in Lip_{1}}E[h(X)-h(Y)],$
where $Lip_{L}$, the class of Lipschitz functions from $\Re^{d}$ to $\Re$ with Lipschitz constant $L,$ is defined as
$Lip_{L}=\{h:|h(x)-h(y)|\leq L||x-y||\}.$
For later reference, we use the notation $||A||_{op}$ to denote the operator norm of a matrix $A,$ i.e.,
$||A||_{op}=\sup_{x:||x||=1}||Ax||.$
The following results from (Gallouët et al. 2018, Propositions 2.2 and 2.3) and (Fang et al. 2019, Theorem 3.1) will be useful to us and so we present them together as a lemma for future reference.
Lemma 1.
Let $g\in Lip_{L}$ and $f$ be the solution to Stein’s equation:
$g(u)-E(g(Z))=\Delta f(u)-u^{T}\nabla f(u),$ (2)
where $\Delta$ and $\nabla$ are the Laplacian and gradient operators, respectively, and $Z\sim N(0,I)$. Then, $f$ has the following regularity properties for any $\beta\in(0,1)$:
  1. 1.
    $||\nabla^{2}f(x)-\nabla^{2}f(y)||_{op}\leq\tilde{C}_{1}(d,\beta)||x-y||^{\beta}L,$ where $\nabla^{2}$ is the Hessian operator.
  2. 2.
    $|\Delta f(x)-\Delta f(y)|\leq\tilde{C_{2}}(d,\beta)||x-y||^{\beta}L$, where
    $\tilde{C}_{1}=C_{1}(d)+\frac{2}{1-\beta},\quad\tilde{C}_{2}=C_{2}(d)+\frac{2d}{1-\beta},\quad C_{1}=2^{3/2}\frac{1+2d\Gamma((1+d)/2)}{d\Gamma(d/2)},\quad C_{2}=2\sqrt{\frac{2d}{\pi}}.$
  3. 3.
    $\langle\nabla^{2}f(x),u_{1}u_{2}^{T}\rangle_{HS}\leq CL||u_{1}||\cdot||u_{2}||\,\forall x,u_{1},u_{2}\in\Re^{d},$
    where $C$ is a constant (independent of $d$ and $\beta$) and $\langle\cdot,\cdot\rangle_{HS}$ is the Hilbert-Schmidt inner product (also called the Frobenius inner product), i.e., the sum of the element-by-element product of two matrices.
$\diamond$
Next, we state a well-known result regarding Ornstein-Uhlenbeck processes which will be useful later to establish the nonasymptotic version of the martingale central limit theorem (for example, the result can be easily inferred from the discussions on generators and linear stochastic differential equations in (Arnold 1974)).
Lemma 2.
Consider the O-U process
$dX_{t}=AX_{t}dt+Bdw_{t},$
where $w$ is the standard $d$-dimensional Wiener process, and $A$ and $B$ are $d\times d$ matrices. Then, the generator of the process $\mathcal{A}$ is the operator given by
$\mathcal{A}f(x):=\lim_{\delta\rightarrow 0}E\left(\frac{f(X_{t+\delta})-f(X_{t})|X_{t}=x}{\delta}\right)=\nabla^{T}f(x)Ax+\frac{1}{2}Tr(\nabla^{2}f(x)BB^{T}),$
where $f:\Re^{d}\rightarrow\Re$ is any twice differentiable function. If the eigenvalues of $A$ have strictly negative real parts, then the stationary distribution of the O-U process is Gaussian with mean zero and covariance $\Sigma,$ which satisfies the Lyapunov equation
$\Sigma A^{T}+A\Sigma+BB^{T}=0.$
Thus, if $\tilde{Z}_{k}\sim N(0,\Sigma),$ then $E(\mathcal{A}f(\tilde{Z}_{k}))=0$ for any twice differentiable function $f.$diamond
It is worth noting that Stein’s equation (2) can be motivated by the above lemma when $A=-I$ and $BB^{T}=2I,$ in which case the right-hand side of (2) is the same as $\mathcal{A}f(u)$. Thus, if $u$ in (2) is replaced by a random vector, then the left-hand side of (2) is a measure of the difference between the distribution of that random vector and a standard Gaussian distribution while the right-hand side will hopefully be small if the distribution of the random vector is close to a standard Gaussian distribution (see, for example, (Barbour 1990, Gotze 1991, Chen et al. 2010)).
2.2 Rate of Convergence
We now state and prove the main result of this section.
Theorem 1.
Let $Z\sim N(0,I)$ and $W_{n}$ be defined as in (1). Then, under Assumption 1, we have
$d_{\mathcal{W}}(W_{n},\Sigma_{\infty}^{1/2}Z)$
$\leq\frac{1}{\sqrt{n}}\sum_{k=1}^{n}\Biggl{[}\frac{\tilde{C}_{1}(d,\beta)||\Sigma_{\infty}^{1/2}||_{op}}{(n-k+1)^{(1+\beta)/2}}E\biggl{(}||\Sigma_{\infty}^{-1/2}m_{k}||^{2+\beta}\biggr{)}+\frac{\tilde{C}_{2}(d,\beta)||\Sigma_{\infty}^{1/2}||_{op}}{(n-k+1)^{(1+\beta)/2}}E\biggl{(}||\Sigma_{\infty}^{-1/2}m_{k}||^{\beta}\biggr{)}$
$\quad-\frac{1}{n-k+1}Tr\Bigl{(}A_{k}\Bigl{(}\Sigma_{\infty}^{-1/2}E(\Sigma_{k})\Sigma_{\infty}^{-1/2}-I\Bigr{)}\Bigr{)}\Biggr{]},$
where $A_{k}$ is a matrix with the property
$||A_{k}||_{op}\leq C\sqrt{n-k+1}||\Sigma_{\infty}^{1/2}||_{op},$
and $C,\tilde{C}_{1},\tilde{C}_{2}$ are the constants in statement (3) of Lemma 1.
Proof.
We first note that
$d_{\mathcal{W}}(W_{n},\Sigma_{\infty}^{1/2}Z)=\sup_{h\in Lip_{1}}E[h(W_{n})-h(\Sigma_{\infty}^{1/2}Z)]=\sup_{h\in Lip_{1}}\frac{1}{\sqrt{n}}E[h(S_{n})-h(T_{1})],$ (3)
where $T_{k}=\Sigma_{\infty}^{1/2}(Z_{k}+\ldots+Z_{n}),$ where $Z_{i}$ are i.i.d. $N(0,I)$ random vectors, independent of the martingale difference sequence. Then, using the Lindeberg decomposition (Röllin 2018), we have
$E\Bigl{(}h(S_{n})-h(T_{1})\Bigr{)}=\sum_{k=1}^{n}E\Bigl{(}h(S_{k}+T_{k+1})-h(S_{k-1}+T_{k})\Bigr{)},$ (4)
where we have defined $T_{n+1}=S_{0}=0.$
The rest of the proof uses the vector extensions of the techniques used in (Röllin 2018) for scalar random variables. As in well known in the Stein’s method literature, going from scalar-valued random variables to random vectors is non-trivial. To deal with random vectors, we use the ideas in (Gallouët et al. 2018, Theorem 3.1) which were developed for i.i.d. random variables. One of our contributions is to show that these ideas can be used for martingales using the Lindeberg decomposition. While doing so, we also have to deal with the asymptotic covariance $\Sigma_{\infty},$ which does not appear in (Röllin 2018, Anastasiou et al. 2019) due to their assumptions, but will be important to us when we study Markov chains.
Motivated by the form of each term in the Lindeberg decomposition, we consider the difference
$h(x)-E(h(a+QZ)),$
where $x,a\in\Re^{d}$ and $Q\in\Re^{d\times d}$ is a symmetric positive definite matrix. The idea is to study this expression and then later, after appropriate conditioning, substitute
$x=S_{k}+T_{k+1},\quad a=S_{k-1},\quad\mathrm{and}\quad Q=\sqrt{n-k+1}\Sigma_{\infty}^{1/2}.$
Note that the choice of $Q$ is motivated by the fact that $T_{k}$ has the same distribution as
$\sqrt{n-k+1}\Sigma_{\infty}^{1/2}Z.$
Next define the function $\tilde{h}$ by $\tilde{h}(u)=h(a+Qu).$ Since $h\in Lip_{1},$ $\tilde{h}\in Lip_{L}$ with $L=||Q||_{op}.$ Further, from Lemma 1, we know that the solution $f$ to the Stein equation
$\tilde{h}(u)-E(\tilde{h}(Z))=\Delta f(u)-u^{T}\nabla f(u)$
has nice regularity properties. Note that Stein’s equation can be rewritten in terms of $h$ as follows:
$h(x)-E(h(a+QZ))$ $=$ $\tilde{h}(Q^{-1}(x-a))-E(\tilde{h}(Z))$
$=$ $\Delta f(Q^{-1}(x-a))-(Q^{-1}(x-a))^{T}\nabla f(Q^{-1}(x-a)).$
Thus,
$E\Bigl{(}h(S_{k}+T_{k+1})-h(S_{k-1}+T_{k})\Bigr{|}\mathcal{F}_{k-1}\Bigr{)}$
$=E\biggl{(}\Delta f\Bigl{(}\frac{\Sigma_{\infty}^{-1/2}}{\sqrt{n-k+1}}(m_{k}+T_{k+1})\Bigr{)}$
$-(m_{k}+T_{k+1})^{T}\frac{\Sigma_{\infty}^{-1/2}}{\sqrt{n-k+1}}\nabla f\Bigl{(}\frac{\Sigma_{\infty}^{-1/2}}{\sqrt{n-k+1}}(m_{k}+T_{k+1})\Bigr{)}|\mathcal{F}_{k-1}\biggr{)}$
$=E\biggl{(}\Delta f\Bigl{(}\frac{\Sigma_{\infty}^{-1/2}}{\sqrt{n-k+1}}m_{k}+\tilde{Z}_{k}\Bigr{)}-(\frac{\Sigma_{\infty}^{-1/2}}{\sqrt{n-k+1}}m_{k}+\tilde{Z}_{k})^{T}\nabla f\Bigl{(}\frac{\Sigma_{\infty}^{-1/2}}{\sqrt{n-k+1}}m_{k}+\tilde{Z}_{k}\Bigr{)}|\mathcal{F}_{k-1}\biggr{)},$
where $\tilde{Z}_{k}\sim N(0,\sqrt{\frac{n-k}{n-k+1}}I)$ has the same distribution as $\frac{1}{\sqrt{n-k+1}}\Sigma_{\infty}^{-1/2}T_{k+1}$ and is independent of all other random variables . Next, we rewrite the expression as a sum of four terms as follows:
$E\Bigl{(}h(S_{k}+T_{k+1})-h(S_{k-1}+T_{k})\Bigr{|}\mathcal{F}_{k-1}\Bigr{)}$
$=E\biggl{(}\frac{n-k}{n-k+1}\Delta f\Bigl{(}\frac{\Sigma_{\infty}^{-1/2}}{\sqrt{n-k+1}}m_{k}+\tilde{Z}_{k}\Bigr{)}-\tilde{Z}_{k}^{T}\nabla f\Bigl{(}\frac{\Sigma_{\infty}^{-1/2}}{\sqrt{n-k+1}}m_{k}+\tilde{Z}_{k}\Bigr{)}|\mathcal{F}_{k-1}\biggr{)}$
$+E\biggl{(}\frac{1}{n-k+1}\Delta f\Bigl{(}\frac{\Sigma_{\infty}^{-1/2}}{\sqrt{n-k+1}}m_{k}+\tilde{Z}_{k}\Bigr{)}|\mathcal{F}_{k-1}\biggr{)}$
$-E\Bigl{(}(\frac{\Sigma_{\infty}^{-1/2}}{\sqrt{n-k+1}}m_{k})^{T}\nabla f(\tilde{Z}_{k})|\mathcal{F}_{k-1}\Bigr{)}$
$-E\biggl{(}(\frac{\Sigma_{\infty}^{-1/2}}{\sqrt{n-k+1}}m_{k})^{T}\Bigl{(}\nabla f\Bigl{(}\frac{\Sigma_{\infty}^{-1/2}}{\sqrt{n-k+1}}m_{k}+\tilde{Z}_{k}\Bigr{)}-\nabla f(\tilde{Z}_{k})\Bigr{)}|\mathcal{F}_{k-1}\biggr{)}$
The first term on the right-hand side of the previous equation is zero by applying Lemma 2 with
$A=-I,\qquad BB^{T}=2\frac{n-k}{n-k+1}I,$
and using the fact that $\Delta f=Tr(\nabla^{2}f).$ The third term is zero by the martingale difference property. Next, let $\Theta\sim Uniform[0,1]$ be a random variable, independent of all other random variables, and using the fundamental theorem of calculus, we have
$E\Bigl{(}h(S_{k}+T_{k+1})-h(S_{k-1}+T_{k})\Bigr{|}\mathcal{F}_{k-1}\Bigr{)}$
$=E\biggl{(}\frac{1}{n-k+1}\Delta f\Bigl{(}\frac{\Sigma_{\infty}^{-1/2}}{\sqrt{n-k+1}}m_{k}+\tilde{Z}_{k}\Bigr{)}|\mathcal{F}_{k-1}\biggr{)}$
$-E\biggl{(}(\frac{\Sigma_{\infty}^{-1/2}}{\sqrt{n-k+1}}m_{k})^{T}\Bigl{[}\nabla^{2}f\Bigl{(}\Theta\frac{\Sigma_{\infty}^{-1/2}}{\sqrt{n-k+1}}m_{k}+\tilde{Z}_{k}\Bigr{)}-\nabla^{2}f(\tilde{Z}_{k})\Bigr{]}(\frac{\Sigma_{\infty}^{-1/2}}{\sqrt{n-k+1}}m_{k})|\mathcal{F}_{k-1}\biggr{)}$
$-E\biggl{(}(\frac{\Sigma_{\infty}^{-1/2}}{\sqrt{n-k+1}}m_{k})^{T}\nabla^{2}f(\tilde{Z}_{k})(\frac{\Sigma_{\infty}^{-1/2}}{\sqrt{n-k+1}}m_{k})|\mathcal{F}_{k-1}\biggr{)}.$
Next, we perform the following sequence of manipulations: first, we apply statement (1) in Lemma 1 to the second term above, use the definition of $\Sigma_{k}$ (in Assumption 1) in the third term, use the fact that $\Delta f=Tr(\nabla^{2}f),$ and finally apply statement (2) in Lemma 1, while noting that $|\Theta|\leq 1$ a.s., to get
$E\Bigl{(}h(S_{k}+T_{k+1})-h(S_{k-1}+T_{k})\Bigr{|}\mathcal{F}_{k-1}\Bigr{)}$
$\leq E\biggl{(}\frac{1}{n-k+1}\Delta f\Bigl{(}\frac{\Sigma_{\infty}^{-1/2}}{\sqrt{n-k+1}}m_{k}+\tilde{Z}_{k}\Bigr{)}|\mathcal{F}_{k-1}\biggr{)}$
$-\frac{1}{n-k+1}Tr\Bigl{(}E(\nabla^{2}f(\tilde{Z}_{k}))\Sigma_{\infty}^{-1/2}\Sigma_{k}\Sigma_{\infty}^{-1/2}\Bigr{)}$
$\qquad+\frac{\tilde{C}_{1}(d,\beta)||\Sigma_{\infty}^{1/2}||_{op}}{(n-k+1)^{(1+\beta)/2}}E\biggl{(}||\Sigma_{\infty}^{-1/2}m_{k}||^{2+\beta}|\mathcal{F}_{k-1}\biggr{)}$
$=E\biggl{(}\frac{1}{n-k+1}\Bigl{(}\Delta f\Bigl{(}\frac{\Sigma_{\infty}^{-1/2}}{\sqrt{n-k+1}}m_{k}+\tilde{Z}_{k}\Bigr{)}-\Delta f(\tilde{Z}_{k})\Bigr{)}|\mathcal{F}_{k-1}\biggr{)}$
$\quad-\frac{1}{n-k+1}Tr\Bigl{(}\Sigma_{\infty}^{-1/2}\Bigl{(}\Sigma_{k}-\Sigma_{\infty}\Bigr{)}\Sigma_{\infty}^{-1/2}E(\nabla^{2}f(\tilde{Z}_{k}))\Bigr{)}$
$+\frac{\tilde{C}_{1}(d,\beta)||\Sigma_{\infty}^{1/2}||_{op}}{(n-k+1)^{(1+\beta)/2}}E\biggl{(}||\Sigma_{\infty}^{-1/2}m_{k}||^{2+\beta}|\mathcal{F}_{k-1}\biggr{)}$
$=\frac{\tilde{C}_{2}(d,\beta)||\Sigma_{\infty}^{1/2}||_{op}}{(n-k+1)^{(1+\beta)/2}}E\biggl{(}||\Sigma_{\infty}^{-1/2}m_{k}||^{\beta}|\mathcal{F}_{k-1}\biggr{)}$
$\quad-\frac{1}{n-k+1}Tr\Bigl{(}\Sigma_{\infty}^{-1/2}\Bigl{(}\Sigma_{k}-\Sigma_{\infty}\Bigr{)}\Sigma_{\infty}^{-1/2}E(\nabla^{2}f(\tilde{Z}_{k}))\Bigr{)}$
$+\frac{\tilde{C}_{1}(d,\beta)||\Sigma_{\infty}^{1/2}||_{op}}{(n-k+1)^{(1+\beta)/2}}E\biggl{(}||\Sigma_{\infty}^{-1/2}m_{k}||^{2+\beta}|\mathcal{F}_{k-1}\biggr{)}.$
We obtain the result by taking an expectation to remove the conditioning, defining $A_{k}$ to be $E(\nabla^{2}f(\tilde{Z})),$ then substituting the above expression in (4) and (3). To obtain the upper bound on $A_{k},$ we use statement (3) in Lemma 1 and the fact that
$||A_{k}||_{op}=\sup_{||u_{1}||=||u_{2}||=1}|\langle A_{k},u_{1}u_{2}^{T}\rangle_{HS}|.$
$\Box$
Note that if $\Sigma_{\infty}$ is such that $\Sigma_{k}\rightarrow\Sigma_{\infty}$ as $k\rightarrow\infty,$ then the rate at which this convergence occurs will determine the rate of convergence in the martingale CLT. One can characterize such a rate explicitly in the case of Markov chains as we will see in the next section.
3 Markov Chain Central Limit Theorem
Consider a time-homogeneous Markov chain $\{X_{k}\}_{k\geq 0}$ on a state space $\mathcal{S}$ with some initial distribution $\eta.$ Let $r(X_{k})\in\Re^{d}$ denote a vector of rewards accrued when in state $X_{k}.$ We are interested in rates of convergence to normality of
$\frac{\sum_{k=1}^{n}r(X_{k})}{\sqrt{n}}.$
As is common in the literature, we will study the above scaled sum by relating the reward process to a martingale via Poisson’s equation (Douc et al. 2018, Makowski and Shwartz 2002, Glynn and Meyn 1996, Glynn and Infanger 2023).
3.1 Finite State Space Markov Chains
In this subsection, we assume that $\mathcal{S}$ is finite and extend the results to more general state spaces in the next subsection.
Assumption 2.
We assume that the Markov chain takes values in finite set $\mathcal{S}$ and is irreducible and aperiodic. $\diamond$
The assumption has a few well-known consequences that will be useful in the proof of our next result (Ross 2013, Asmussen 2003):
  1. 1.
    There exists a unique stationary distribution $\pi$ and the Markov chain is uniformly geometrically ergodic, i.e., there exists $K_{1}>0$ and $\rho\in[0,1)$ such that $\forall x\in\mathcal{S},$ we have
    $||P(X_{n}=y|X_{0}=x)-\pi(y)||_{TV}\leq K_{1}\rho^{n},$
  2. 2.
    There exists a solution $V:S\rightarrow\Re^{d}$ to the following Poisson’s equation:
    $\bar{r}=r(x)+E(V(X_{1})|X_{0}=x)-V(x)\qquad\forall x\in S,$ (5)
    where $\bar{r}=E_{y\sim\pi}(r(y)),$ such that $||V(x)||\leq M_{V},\forall k\geq 0$ for some $M_{V}<\infty.$
  3. 3.
    Define $m_{k}=V(X_{k})-E(V(X_{k})|X_{k-1})$ for $k\geq 1.$ Clearly, $E(m_{k}|\mathcal{F}_{k-1})=0$ and $\{m_{k}\}_{k\geq 1}$ is a martingale difference sequence with respect to the filtration $\{\mathcal{F}_{k}\}_{k\geq 0}$ satisfying Assumption 1 with $\mathcal{F}_{k}=\sigma(X_{0},X_{1},X_{2},\ldots,X_{k})$. Clearly, from the previous bullet item, there exists constants $M_{2+\beta}$ and $M_{\beta}$ such that
    $E(||m_{k}||^{2+\beta})\leq M_{2+\beta}\mbox{ and }E(||m_{k}||^{\beta})\leq M_{\beta}\qquad\forall k\geq 1.$
  4. 4.
    While Theorem 1 holds for any $\Sigma_{\infty}\succ 0$, in this subsection we will find it convenient to define it to be
    $\Sigma_{\infty}=\sum_{i,j\in S}\pi_{i}P_{ij}(V(j)-E(V(X_{1})|X_{0}=i))(V(j)-E(V(X_{1})|X_{0}=i))^{T},$ (6)
    where $P_{ij}$ is the $(i,j)^{\rm th}$ element of the probability transition matrix. Due to the existence of a solution to Poisson’s equation, $\Sigma_{\infty}$ is well-defined.
We now state the main result of this section, and note that the constants hidden in the $\mathcal{O}$-notation can be made explicit by applying the bounds assumed in Assumption 2 to the right-hand side of the bound in Theorem 1 as outlined in the proof of the theorem. However, this would lead to messy expressions.
Theorem 2.
Let
$U_{n}=\frac{1}{\sqrt{n}}\sum_{k=1}^{n}(r(X_{k-1})-\bar{r}).$
Then, under Assumption 2,
  1. 1.
    for any $\beta\in(0,1),$ $d_{\mathcal{W}}(U_{n},\Sigma_{\infty}^{1/2}Z)=\mathcal{O}\Bigl{(}\frac{1}{(1-\beta)n^{\beta/2}}\Bigr{)},$ and
  2. 2.
    $d_{\mathcal{W}}(U_{n},\Sigma_{\infty}^{1/2}Z)=\mathcal{O}\Bigl{(}\frac{\log n}{\sqrt{n}}\Bigr{)},$
where $Z\sim N(0,I).$
Proof.
We start with the following standard use of Poisson’s equation and the definition of $m_{k}$ in Assumption 2 (see, for example, (Douc et al. 2018)) to write
$r(X_{k-1})-\bar{r}$ $=V(X_{k-1})-E(V(X_{k})|X_{k-1})$
$=V(X_{k-1})-V(X_{k})+V(X_{k})-E(V(X_{k})|X_{k-1})$
$=V(X_{k-1})-V(X_{k})+m_{k}$
Thus,
$U_{n}=W_{n}+\frac{V(X_{0})-V(X_{n})}{\sqrt{n}},\quad\mathrm{where}\quad W_{n}=\frac{1}{\sqrt{n}}\sum_{k=1}^{n}m_{k}.$ (7)
For any $h\in Lip_{1},$
$h(U_{n})\leq h(W_{n})+||U_{n}-W_{n}||=h(W_{n})+\frac{2M_{V}}{\sqrt{n}}.$
This implies
$d_{\mathcal{W}}(U_{n},\Sigma_{\infty}^{1/2}Z)\leq d_{\mathcal{W}}(W_{n},\Sigma_{\infty}^{1/2}Z)+\frac{2M_{V}}{\sqrt{n}}.$
We can now apply Theorem 1 to bound $d_{\mathcal{W}}(W_{n},\Sigma_{\infty}^{1/2}Z).$ By Assumption 2, the first two terms in the bound in Theorem 1 go to zero at rate
$\frac{1}{\sqrt{n}}\sum_{k=1}^{n}\frac{1}{(n-k+1)^{(1+\beta)/2}}=\frac{1}{\sqrt{n}}\sum_{j=1}^{n}\frac{1}{j^{(1+\beta)/2}}\leq\frac{K_{2}}{n^{\beta/2}},$
for some universal constant $K_{2}.$ For the third term in the bound in Theorem 1,
$\frac{1}{n-k+1}Tr\Bigl{(}A_{k}\Sigma_{\infty}^{-1/2}\Bigl{(}E(\Sigma_{k})-\Sigma_{\infty}\Bigr{)}\Sigma_{\infty}^{-1/2}\Bigr{)}$ $\leq\frac{1}{n-k+1}||A_{k}||_{HS}||\Sigma_{\infty}^{-1/2}E(\Sigma_{k})\Sigma_{\infty}^{-1/2}-I||_{HS}$
$\leq\frac{1}{n-k+1}\sqrt{d}||A_{k}||_{op}||\Sigma_{\infty}^{-1/2}E(\Sigma_{k})\Sigma_{\infty}^{-1/2}-I||_{HS}$
$\leq C\sqrt{d}||\Sigma_{\infty}^{1/2}||_{op}||\Sigma_{\infty}^{-1/2}E(\Sigma_{k})\Sigma_{\infty}^{-1/2}-I||_{HS},$ (8)
where the last step follows from Theorem 1.
Next, letting $p_{i}(k)=Prob(X_{k}=i),$ we get
$E(\Sigma_{k})$ $=\sum_{i,j\in S}p_{i}(k)P_{ij}(V(j)-E(V(X_{1})|X_{0}=i))(V(j)-E(V(X_{1})|X_{0}=i))^{T}$
$=\sum_{i,j\in S}(p_{i}(k)-\pi_{i})P_{ij}(V(j)-E(V(X_{1})|X_{0}=i))(V(j)-E(V(X_{1})|X_{0}=i))^{T}+\Sigma_{\infty},$
which implies
$||\Sigma_{\infty}^{-1/2}E(\Sigma_{k})\Sigma_{\infty}^{-1/2}-I||_{HS}$
$\leq||\Sigma_{\infty}^{-1/2}||_{HS}^{2}\sum_{i,j\in S}|p_{i}(k)-\pi_{i}|\cdot||P_{ij}(V(j)-E(V(X_{1})|X_{0}=i))(V(j)-E(V(X_{1})|X_{0}=i))^{T}||_{HS}.$
The above expression goes to zero geometrically fast and thus,
$\frac{1}{\sqrt{n}}\sum_{k=1}^{n}\frac{1}{n-k+1}Tr\Bigl{(}A_{k}\Sigma_{\infty}^{-1/2}\Bigl{(}E(\Sigma_{k})-\Sigma_{\infty}\Bigr{)}\Sigma_{\infty}^{-1/2}\Bigr{)}\leq\frac{K_{3}}{\sqrt{n}},$
where $K_{3}$ is a problem-dependent constant.
As in (Gallouët et al. 2018), we can obtain the following corollary by choosing $\beta=1-2/\log n$ and noting that $\tilde{C}_{1}$ and $\tilde{C}_{2}$ become $O(\log n)$ in that case. It is worth noting that if we set $\beta=1-2/\log n,$ we require third moments to exist for the non-asymptotic martingale central limit theorem to hold. This condition is satisfied since the solution to Poisson’s equation is bounded under the assumption that the Markov chain has a finite state space and is irreducible and aperiodic. $\Box$
3.2 Extension to General State-Space Markov Chains
We now extend the results to general state-space Markov chains. The ideas are similar to those of the finite state-space case, but we need additional assumptions and work to prove the result. To avoid unnecessary repetition, we use much of the same notation as in the previous section, while additional terminology and machinery are borrowed from (Meyn and Tweedie 2012). Not surprisingly, we need stronger assumptions than the conditions needed in (Meyn and Tweedie 2012) to establish the CLT since we are interested in obtaining the rate of convergence.
Theorem 3.
Let $X$ be a $\psi$-irreducible, aperiodic Markov chain over a state space $\mathcal{S}$. Suppose there exists a Lyapunov function $\mathscr{V}\geq 1$, with $E(V(X_{0}))<\infty$, such that the following drift condition is satisfied:
$E(\mathscr{V}(X_{k+1})|X_{k})\leq\lambda\mathscr{V}(X_{k})+L$ (9)
for some $\lambda\in[0,1)$ and $L\in[0,\infty).$ (Note that it is well known that one can equivalently assume a traditional Foster-type drift condition and strengthen it to (9) by imposing reasonable conditions on the increments of the Lyapunov function; see (Popov 1977, Hajek 1982, Spieksma and Tweedie 1994) and (Meyn and Tweedie 2012, Theorem 16.3.1).) Then, the following results hold:
  1. 1.
    For any $\beta\in(0,1)$, if $|r(x)|^{2+\beta}\leq\mathscr{V}(x)\,\forall x\in S,$ then
    $d_{\mathcal{W}}(U_{n},\Sigma_{\infty}^{1/2}Z)=\mathcal{O}\Bigl{(}\frac{1}{(1-\beta)n^{\beta/2}}\Bigr{)},$
    where $Z\sim N(0,I).$
  2. 2.
    If $|r(x)|^{3}\leq\mathscr{V}(x)\,\forall x\in S,$ then
    $d_{\mathcal{W}}(U_{n},\Sigma_{\infty}^{1/2}Z)=\mathcal{O}\Bigl{(}\frac{\log n}{\sqrt{n}}\Bigr{)}.$
Proof.
We will only present the additional steps needed beyond the finite-state case to obtain the second statement in the theorem. The proof of the first statement is nearly identical.
Noting the decomposition in (7) and as in the rest of the proof of Theorem 2, we need to prove the following four facts:
  1. 1.
    $E(||m_{k}||^{3}$) is uniformly bounded.
  2. 2.
    $E(||V(X_{0})-V(X_{n})||)$ is bounded uniformly in $n.$
  3. 3.
    $\Sigma_{\infty}=E_{\pi}((V(X_{1})-E(V(X_{1})|X_{0}))(V(X_{1})-E(V(X_{1})|X_{0}))^{T}$ exists, where $E_{\pi}$ denotes that $X_{0}\sim\pi.$
  4. 4.
    $Tr\Bigl{(}A_{k}\Bigl{(}\Sigma_{\infty}^{-1/2}E(\Sigma_{k})\Sigma_{\infty}^{-1/2}-I\Bigr{)}\Bigr{)}$ decays sufficiently fast as a function of $k.$
Once these facts are proved, the proof of the theorem is identical to the proof of Theorem 2.
Proof of (1):
From (9), we have:
$E(\mathscr{V}^{1/3}(X_{k+1})|X_{k})$ $\leq[E(\mathscr{V}(X_{k+1})|X_{k})]^{1/3}$
$\leq(\lambda\mathscr{V}(X_{k})+L)^{1/3}$
$\leq\lambda^{1/3}\mathscr{V}^{1/3}(X_{k})+L^{1/3},$
where the first inequality is Jensen’s inequality, the second is the from the drift condition (9) and the third can be seen by cubing the expressions in the second and third lines. Equivalently,
$E(\mathscr{V}^{1/3}(X_{k+1})|X_{k})-\mathscr{V}^{1/3}(X_{k})\leq-(1-\lambda^{1/3})\mathscr{V}^{1/3}(X_{k})+L^{1/3}.$ (10)
Now using the bound on $|r_{i}(x)|$ and Theorem 17.4.2 (and the discussion prior to it) in (Meyn and Tweedie 2012), we can conclude that there exists a solution to Poisson’s equation which satisfies the following bound:
$|V_{i}(x)|\leq B(\mathscr{V}^{1/3}(x)+1)\Rightarrow|V_{i}(x)|^{3}\leq 8B^{3}(\mathscr{V}(x)+1)\quad\forall x,$
where $V_{i}$ is the $i^{\rm th}$ component of $V$ and we have used the fact $(a+b)^{3}\leq(2\max(a,b))^{3}\leq 8(a^{3}+b^{3})$ for positive $a$ and $b.$
Next, considering the $i^{\rm th}$ component of $m_{k},$ we have
$|m_{ki}|^{3}$ $\leq 8(|E(V_{i}(X_{k}))|^{3}+|E(V_{i}(X_{k})|X_{k-1})|^{3}$
$=8(E(V_{i}^{3}(X_{k}))+E(V_{i}^{3}(X_{k})|X_{k-1}).$
Letting $E(\mathscr{V}(X_{\infty}))$ denote the expectation of $\mathscr{V}$ with respect to the stationary distribution, we have
$E(|m_{ki}|^{3})\leq 16E(V^{3}(X_{k}))$ $\leq 144B^{3}(E(\mathscr{V}(X_{k}))+1)$
$\leq 144B^{3}\left(E(\mathscr{V}(X_{\infty}))+1+|E(\mathscr{V}(X_{k}))-E(\mathscr{V}(X_{\infty}))|\right)$
$\leq 144B^{3}\left(E(\mathscr{V}(X_{\infty}))+1+RE(V(X_{0}))\rho^{k}\right),$
for some $\rho\in[0,1),$ $R\in[0,\infty)$, where the last step follows from the geometric ergodicity theorem (Meyn and Tweedie 2012, Theorem 15.0.1). Furthermore, from the $f$-norm ergodic theorem (Meyn and Tweedie 2012, Theorem 14.0.1), we have $E(V(X_{\infty}))<\infty,$ which proves that $E(||m_{k}||^{3})$ is uniformly bounded in $k.$
Proof of (2):
It suffices to show that $E(||V(X_{n})||)$ is uniformly bounded.
$E(|V_{i}(X_{n})|)\leq BE(\mathscr{V}^{1/3}(X_{n})+1)$ $\leq BE(\mathscr{V}(X_{n})+1)$
$\leq B\Big{(}|E(\mathscr{V}(X_{n})-\mathscr{V}(X_{\infty}))|+|E(\mathscr{V}(X_{\infty}))|+1\Big{)}.$
As discussed above, $|E(\mathscr{V}(X_{n})-\mathscr{V}(X_{\infty}))|$ is uniformly bounded in view of the geometric ergodic theorem, thus concluding the proof.
Proof of (3):
Consider the $(i,j)^{\rm th}$ element of $\Sigma_{\infty}:$
$Cov_{\pi}(E(V_{i}(X_{1}|X_{0}),E(V_{j}(X_{1}|X_{0}))\leq\sqrt{Var_{\pi}(E(V_{i}(X_{1})|X_{0}))Var_{\pi}(E(V_{i}(X_{1})|X_{0}))}.$
Now
$Var_{\pi}(E(V_{i}(X_{1})|X_{0})\leq E_{\pi}(E(V_{i}(X_{1})|X_{0}))^{2}$ $\leq B^{2}E_{\pi}(E(\mathscr{V}^{1/3}(X_{1})|X_{0})+1)^{2}$
$\leq 2B^{2}(E_{\pi}((E(\mathscr{V}^{1/3}(X_{1})|X_{0}))^{2}+1)$
$\leq 2B^{2}(E_{\pi}((E(\mathscr{V}(X_{1})|X_{0}))^{2/3})+1)$
$\leq 2B^{2}(E_{\pi}(E(\mathscr{V}(X_{1})|X_{0}))+1)$
$\leq 2B^{2}(E_{\pi}(\mathscr{V}(X_{1}))+1),$
which is finite, thus proving (3).
Proof of (4):
Consider the absolute value of the expectation of the $(i,j)^{\rm th}$ term of $\Sigma_{k}=E(m_{k}m_{k}^{T}|X_{k-1}):$
$|Cov(E(V_{i}(X_{k})|X_{k-1}),E(V_{j}(X_{k})|X_{k-1})|X_{k})|$ $\leq B^{2}E\Big{(}(E(\mathscr{V}^{1/3}(X_{k})|X_{k-1})+E(\mathscr{V}^{1/3}(X_{k})))^{2}\Big{)}$
$\leq 2B^{2}E\Big{(}(E(\mathscr{V}^{2/3}(X_{k})|X_{k-1})+E(\mathscr{V}^{2/3}(X_{k})))\Big{)}$
$\leq 2B^{2}E\Big{(}E(\mathscr{V}(X_{k})|X_{k-1})+E(\mathscr{V}(X_{k}))\Big{)}$
$=4B^{2}E(\mathscr{V}(X_{k}))$
$=4B^{2}\Big{(}|E(\mathscr{V}(X_{k}))-E(\mathscr{V}(X_{\infty}))|+|E(\mathscr{V}(X_{\infty}))|\Big{)}$
Thus, the absolute value of the $(i,j)^{\rm th}$ term of $\Sigma_{k}-\Sigma_{\infty}$ is upper bounded by a constant by geometric ergodicity. $\Box$
4 An Application to TD Learning
Temporal Difference (TD) Learning is a common method to learn the performance of a policy in reinforcement learning; see (Bertsekas and Tsitsiklis 1996, Sutton and Barto 2018) for details. In (Tsitsiklis and Van Roy 1996), it was shown that TD learning can be represented as a jump Markovian linear dynamical system as follows:
$\theta_{k+1}=\theta_{k}-\epsilon_{k}(A(X_{k})\theta_{k}+b(X_{k})),$ (11)
where $\theta_{k}\in\Re^{d},$ $\epsilon_{k}$ is the stepsize, $\{X_{k}\}_{k\geq 0}$ is a finite state-space Markov chain, $A(X_{k})$ is a $d\times d$ matrix, $b(X_{k})$ is a $d$-dimensional vector and $\theta_{0}$ being some constant vector. We make the following assumptions.
Assumption 3.
We assume that the TD learning algorithms satisfies the following conditions:
  1. 1.
    $\{X_{k}\}$ is an irreducible, aperiodic, finite state Markov chain. So it satisfies Assumption 2, item (1).
  2. 2.
    Let $\bar{A}:=E_{X_{k}\sim\pi}(A(X_{k})).$ Then, $-\bar{A}$ is a Hurwitz matrix, i.e., all eigenvalues have strictly negative real parts.
$\diamond$
The first item in the assumption is standard in the reinforcement learning literature (Sutton and Barto 2018), while the second item has been established for TD learning in (Tsitsiklis and Van Roy 1996).
Define $\theta^{*}$ by $\bar{A}\theta^{*}+\bar{b}=0,$ where $\bar{b}=E_{X_{k}\sim\pi}(b(X_{k})).$ Here, we are interested in the Polyak-Ruppert averaged version of the TD-learning algorithm, i.e.,
$\bar{\theta}_{n}=\frac{1}{n}\sum_{k=1}^{n}\theta_{k}.$
Our main result of this section characterizes the rate of convergence of this algorithm for an appropriately chosen $\{\epsilon_{k}\}$. As in the previous theorem, we have chosen to write the result using the $\mathcal{O}$ notation rather than making the constants explicit, but these constants can be made explicit by following the details the proof of the theorem.
Theorem 4.
Under Assumption 3, and with $\epsilon_{k}=1/(k+1)^{\delta},$ $\delta\in(0.5,1)$ the following rate of convergence holds:
$d_{\mathcal{W}}(\sqrt{n}(\bar{\theta}_{n}-\theta^{*}),(\bar{A}^{-1}\Sigma_{\infty}\bar{A}^{-T})^{1/2}Z)=\mathcal{O}\Bigl{(}\max\left(\frac{\sqrt{\log n}}{n^{\delta-0.5}},\frac{1}{n^{(1-\delta)/2}}\right)\Bigr{)},$
where $\Sigma_{\infty}$ has been defined in Assumption 2, with $r(X_{k})$ as in Assumption 3.
Proof.
Let $\Delta_{k}=\theta_{k}-\theta^{*}.$ It follows from (11) that
$\Delta_{k+1}=$ $(I-\epsilon_{k}\bar{A})\Delta_{k}-\epsilon_{k}(A(X_{k})-\bar{A})\Delta_{k}-\epsilon_{k}\Bigl{(}(A(X_{k})-\bar{A})\theta^{*}+(b(X_{k})-\bar{b})\Bigr{)}.$
Thus,
$\sqrt{n}(\bar{\theta}_{n}-\theta^{*})$ $=\frac{1}{\sqrt{n}}\sum_{k=1}^{n}\prod_{l=0}^{k-1}(I-\epsilon_{l}\bar{A})\Delta_{0}$
$-\frac{1}{\sqrt{n}}\sum_{k=1}^{n}\sum_{j=0}^{k-1}\prod_{l=j+1}^{k-1}(I-\epsilon_{l}\bar{A})\epsilon_{j}(A(X_{j})-\bar{A})\Delta_{j}$
$-\frac{1}{\sqrt{n}}\sum_{k=1}^{n}\sum_{j=0}^{k-1}\prod_{l=j+1}^{k-1}(I-\epsilon_{l}\bar{A})\epsilon_{j}\Bigl{(}(A(X_{j})-\bar{A})\theta^{*}+(b(X_{j})-\bar{b})\Bigr{)}$
$=\frac{1}{\sqrt{n}}\sum_{k=1}^{n}\prod_{l=0}^{k-1}(I-\epsilon_{l}\bar{A})\Delta_{0}$
$-\frac{1}{\sqrt{n}}\sum_{j=0}^{n-1}\epsilon_{j}\sum_{k=j+1}^{n}\prod_{l=j+1}^{k-1}(I-\epsilon_{l}\bar{A})(A(X_{j})-\bar{A})\Delta_{j}$
$-\frac{1}{\sqrt{n}}\sum_{j=0}^{n-1}\epsilon_{j}\sum_{k=j+1}^{n}\prod_{l=j+1}^{k-1}(I-\epsilon_{l}\bar{A})\Bigl{(}(A(X_{j})-\bar{A})\theta^{*}+(b(X_{j})-\bar{b})\Bigr{)},$
where we have used the convention that any product of matrices with the lower index greater than the upper index is equal to the identity matrix. Next, we define
$\Upsilon_{j}^{n}=\epsilon_{j}\sum_{k=j+1}^{n}\prod_{l=j+1}^{k-1}(I-\epsilon_{l}\bar{A})-\bar{A}^{-1}\quad\mbox{ and }\quad\Phi_{j}^{n}=\epsilon_{j}\sum_{k=j}^{n-1}\prod_{l=j}^{k-1}(I-\epsilon_{l}\bar{A})-\bar{A}^{-1}.$ (12)
Properties of $\Phi_{j}^{n}$ are discussed in (Polyak and Juditsky 1992) and $\Upsilon_{j}^{n}=\frac{\epsilon_{j}}{\epsilon_{j+1}}\Phi_{j}^{n+1}.$ Then, we have
$\sqrt{n}(\bar{\theta}_{n}-\theta^{*})$ $=\frac{1}{\sqrt{n}}\sum_{k=1}^{n}\prod_{l=0}^{k-1}(I-\epsilon_{l}\bar{A})\Delta_{0}$ (13)
$-\frac{1}{\sqrt{n}}\sum_{j=0}^{n-1}\bar{A}^{-1}(A(X_{j})-\bar{A})\Delta_{j}$ (14)
$-\frac{1}{\sqrt{n}}\sum_{j=0}^{n-1}\Upsilon_{j}^{n}(A(X_{j})-\bar{A})\Delta_{j}$ (15)
$-\frac{1}{\sqrt{n}}\sum_{j=0}^{n-1}\Upsilon_{j}^{n}\Bigl{(}(A(X_{j})-\bar{A})\theta^{*}+(b(X_{j})-\bar{b})\Bigr{)}$ (16)
$-\frac{1}{\sqrt{n}}\sum_{j=0}^{n-1}\bar{A}^{-1}\Bigl{(}(A(X_{j})-\bar{A})\theta^{*}+(b(X_{j})-\bar{b})\Bigr{)}$ (17)
Now, we will complete the proof in several steps by considering each of the terms above.
Step 1: Equation (13).
Since
$\left|\left|\sum_{k=1}^{n}\prod_{l=0}^{k-1}(I-\epsilon_{l}\bar{A})\epsilon_{0}\right|\right|\leq||\Upsilon_{0}^{n}||_{op}+||\bar{A}^{-1}||_{op},$ (18)
and, from (Polyak and Juditsky 1992), $\Phi_{0}^{n}$ (and hence, $\Upsilon_{0}^{n}$) is bounded independent of $n$, we conclude that this term is $\mathcal{O}(\frac{1}{\sqrt{n}}).$
Step 2: Expression (14).
To analyze (14), we consider the solution $V_{A}$ to the following matrix-valued Poisson’s equation:
$A(X_{k-1})-\bar{A}=V_{A}(X_{k-1})-E(V_{A}(X_{k})|X_{k-1}).$
Let the associated martingale difference sequence $\mu$ be defined by
$\mu_{k}=V_{A}(X_{k})-E(V_{A}(X_{k})|X_{k-1}).$
Thus, up to a sign change, (14) can be written as
$=\frac{1}{\sqrt{n}}\sum_{j=0}^{n-1}\bar{A}^{-1}(V_{A}(X_{j})-V_{A}(X_{j+1}))\Delta_{j}+\frac{1}{\sqrt{n}}\sum_{j=0}^{n-1}\bar{A}^{-1}\mu_{j+1}\Delta_{j}$
$=\frac{1}{\sqrt{n}}\sum_{j=0}^{n-1}\bar{A}^{-1}(V_{A}(X_{j})\Delta_{j}-V_{A}(X_{j+1})\Delta_{j+1})$ (19)
$+\frac{1}{\sqrt{n}}\sum_{j=0}^{n-1}\bar{A}^{-1}V_{A}(X_{j+1})(\Delta_{j+1}-\Delta_{j})$ (20)
$+\frac{1}{\sqrt{n}}\sum_{j=0}^{n-1}\bar{A}^{-1}\mu_{j+1}\Delta_{j}$ (21)
After telescoping,
[…]
where in the last step we have used the fact that solution to the matrix Poisson solution is bounded due to the assumptions on our Markov chain and $E(||\Delta_{n}||^{2})$ is bounded from (Srikant and Ying 2019). Next, due to the fact that our Markov chain’s state space is finite, there exists constants $\kappa_{3}$ and $\kappa_{4}$ such that
[…]
where we have again used the fact $E(||\Delta_{n}||^{2})$ is uniformly bounded. To study (21), we use the martingale difference property of $\mu$ and the fact that it is bounded to conclude that there exists a constant $\kappa_{5}$ such that
[…]
Step 3: Expression (15).
We write (15) as
$\frac{1}{\sqrt{n}}\sum_{j=0}^{n-1}\Upsilon_{j}^{n}(A(X_{j})-\bar{A})\Delta_{j}$ $=\frac{1}{\sqrt{n}}\sum_{j=0}^{n-1}\Upsilon_{j}^{n}(V_{A}(X_{j})-V_{A}(X_{j+1})+\mu_{j+1})\Delta_{j}$ (22)
$=\frac{1}{\sqrt{n}}\sum_{j=0}^{n-1}(\Upsilon_{j}^{n}V_{A}(X_{j})\Delta_{j}-\Upsilon_{j+1}^{n}V_{A}(X_{j+1})\Delta_{j+1})$ (23)
$+\frac{1}{\sqrt{n}}\sum_{j=0}^{n-1}\left(\Upsilon_{j+1}^{n}V_{A}(X_{j+1})(\Delta_{j+1}-\Delta_{j})\right)$ (24)
$+\frac{1}{\sqrt{n}}\sum_{j=0}^{n-1}\left((\Upsilon_{j+1}^{n}-\Upsilon_{j}^{n})V_{A}(X_{j+1})\Delta_{j}\right)$ (25)
$+\frac{1}{\sqrt{n}}\sum_{j=0}^{n-1}\Upsilon_{j}^{n}\mu_{j+1}\Delta_{j}$ (26)
[…]
We can write $\Upsilon_{j+1}^{n}-\Upsilon_{j}^{n}$ as follows:
$\epsilon_{j+1}\sum_{k=j+2}^{n}\prod_{l=j+2}^{k-1}(I-\epsilon_{l}\bar{A})-\epsilon_{j}\sum_{k=j+1}^{n}\prod_{l=j+1}^{k-1}(I-\epsilon_{l}\bar{A})$
$=\epsilon_{j+1}\sum_{k=j+2}^{n}\prod_{l=j+2}^{k-1}(I-\epsilon_{l}\bar{A})-\epsilon_{j+1}\sum_{k=j+1}^{n}\prod_{l=j+1}^{k-1}(I-\epsilon_{l}\bar{A})+(\epsilon_{j+1}-\epsilon_{j})\sum_{k=j+1}^{n}\prod_{l=j+1}^{k-1}(I-\epsilon_{l}\bar{A})$
$=\epsilon_{j+1}\sum_{k=j+2}^{n}\prod_{l=j+2}^{k-1}(I-\epsilon_{l}\bar{A})-\epsilon_{j+1}\sum_{k=j+2}^{n}\prod_{l=j+1}^{k-1}(I-\epsilon_{l}\bar{A})$
$+\epsilon_{j+1}\sum_{k=j+2}^{n}\prod_{l=j+1}^{k-1}(I-\epsilon_{l}\bar{A})-\epsilon_{j+1}\sum_{k=j+1}^{n}\prod_{l=j+1}^{k-1}(I-\epsilon_{l}\bar{A})+(\epsilon_{j+1}-\epsilon_{j})\sum_{k=j+1}^{n}\prod_{l=j+1}^{k-1}(I-\epsilon_{l}\bar{A})$
$=\epsilon_{j+1}^{2}\sum_{k=j+2}^{n}\prod_{l=j+2}^{k-1}(I-\epsilon_{l}\bar{A})\bar{A}-\epsilon_{j+1}(1-\epsilon_{j}\bar{A})+\frac{\epsilon_{j+1}-\epsilon_{j}}{\epsilon_{j}}(\Upsilon_{j}^{n}+\bar{A}^{-1})$
$=\epsilon_{j+1}(\Upsilon_{j+1}^{n}+\bar{A}^{-1})\bar{A}-\epsilon_{j+1}I+\frac{\epsilon_{j+1}-\epsilon_{j}}{\epsilon_{j}}(\Upsilon_{j}^{n}+\bar{A}^{-1}).$
[…]
Step 4: Expression (16).
We note that with $r(X_{k})=A(X_{k})\theta^{*}+b(X_{k}),$ the solution of Poisson’s equation (5) satisfies the conditions in items (2)-(4) of Assumption 2 because our Markov chain is finite-state, irreducible and aperiodic. Using the notation in items (2)-(3) of Assumption 2, (16) can be written as
$\frac{1}{\sqrt{n}}\sum_{j=0}^{n-1}\Upsilon_{j}^{n}\Bigl{(}(A(X_{j})-\bar{A})\theta^{*}+(b(X_{j})-\bar{b})\Bigr{)}$ $=\frac{1}{\sqrt{n}}\sum_{j=0}^{n-1}\Upsilon_{j}^{n}\Bigl{(}V(X_{j})-V(X_{j+1})+m_{j+1}\Bigr{)}$
$=\frac{1}{\sqrt{n}}\sum_{j=0}^{n-1}(\Upsilon_{j}^{n}V(X_{j})-\Upsilon_{j+1}^{n}V(X_{j+1}))$
$+\frac{1}{\sqrt{n}}\sum_{j=0}^{n-1}(\Upsilon_{j+1}^{n}-\Upsilon_{j}^{n})V(X_{j+1})$
$+\frac{1}{\sqrt{n}}\sum_{j=0}^{n-1}\Upsilon_{j}^{n}m_{j+1}$ (27)
The uniform boundedness of $\Upsilon_{j}^{n}$ and $V(x)$ proves that the first two terms after the last equality above go to zero at rate $\mathcal{O}(\frac{1}{\sqrt{n}}).$ Next, using the fact that $\{m_{j}\}$ is a bounded martingale difference sequence, we get
$E(\left(||\frac{1}{\sqrt{n}}\sum_{j=0}^{n-1}\Upsilon_{j}^{n}m_{j+1}||^{2}\right)=\frac{1}{n}E\left(\sum_{j=0}^{n-1}||\Upsilon_{j}^{n}m_{j+1}||^{2}\right)\leq\kappa_{6}\frac{1}{n}\sum_{j=0}^{n-1}||\Upsilon_{j}^{n}||^{2}$
for some constant $\kappa_{6}.$ It is known from (Polyak and Juditsky 1992) that this term goes to zero as $n\rightarrow\infty$. In Appendix B, we show that this term is $\mathcal{O}(1/n^{1-\delta}).$
Conclusion of the Proof:
To finally establish the result stated in the theorem, we observe that, for random variables $\Xi_{i},$ $i=1,2,3,4,5,$ and $h\in Lip_{1},$
$E\Big{(}h(\sum_{i=1}^{5}\Xi_{i})-(h((\bar{A}^{-1}\Sigma_{\infty}\bar{A}^{-T})^{1/2}Z)\Big{)}$
$=\sum_{i=1}^{4}E\left(h(\sum_{k=i}^{5}\Xi_{k})-h(\sum_{j=i+1}^{5}\Xi_{j})\right)+E(h(\Xi_{5})-h((\bar{A}^{-1}\Sigma_{\infty}\bar{A}^{-T})^{1/2}Z))$
$\leq\sum_{i=1}^{4}E(||\Xi_{i}||)+E(h(\Xi_{5}))-h((\bar{A}^{-1}\Sigma_{\infty}\bar{A}^{-T})^{1/2}Z)$
Let $\Xi_{i},$ $i=1,2,3,4$ be the terms in (13)-(16) and $\Xi_{5}$ be (17). The different steps of the proof have shown that each of $E(||\Xi_{i}||)$ for $i=1,2,3,4$ is either $\mathcal{O}(\frac{\sqrt{\log n}}{n^{(\delta-0.5})/2})$ or $\mathcal{O}(\frac{1}{n^{1-\delta}})$. Now we apply Theorem 2 to $\Xi_{5}.$ Since $\Xi_{5}$ converges to a Gaussian distribution in the Wasserstein distance at rate $\mathcal{O}(1/n^{\beta})$ and $\beta<1$ can be arbitrary due to our assumptions on the Markov chain, we have proved the theorem. $\Box$
5 Conclusions
We derived a non-asymptotic central limit theorem for martingales and Markov chains, and applied the results to get rate of convergence in the central limit theorem for TD learning with averaging. The results for martingales and Markov chains may be more broadly applicable to study other reinforcement learning algorithms as well as other algorithms such as nonlinear stochastic approximation and stochastic gradient descent.
There are several possible avenues for further work:
  • One can try to use the martingale CLT results in this paper to strengthen the results and allow for Markovian noise in (Anastasiou et al. 2019). However, these extensions would require us to go beyond the linear setting in this paper and consider nonlinear problems. Along similar lines, it would be also interesting to see if our results can be used to study contractive stochastic approximation as in (Chen et al. 2010) or other stochastic approximation schemes (Borkar 2009, Kushner and Yin 2003).
  • Under some conditions, the rate of convergence in the central limit theorem for i.i.d. random vectors is $\mathcal{O}(1/\sqrt{n})$ (Fang et al. 2019). It would be interesting to see if the $\mathcal{O}(\log n/\sqrt{n})$ rate of convergence in Theorem 1 can be improved to $\mathcal{O}(1/\sqrt{n})$ under possibly additional conditions than the ones used in this paper. The Edgeworth expansion for scalar functions of Markov chains in (Kontoyiannis and Meyn 2003) suggests that the $\log n$ could possibly be eliminated for Markov chains as well. The reason for $\log n$ in our result is due to the fact that we use the regularity results for the solution to Stein’s equation in (Gallouët et al. 2018) after we do the Lindeberg decomposition in the proof of the rate of convergence of martingale CLT. A different approach may be required to eliminate the $\mathcal{O}(\log n)$ term in the numerator.
  • The approach in (Gallouët et al. 2018) that we have used is known to be not tight in terms of the dependence on the dimension $d$ for i.i.d. random vectors. Under some conditions, the dependence on $d$ can be considerably strengthened in the i.i.d. case (Courtade et al. 2019, Zhai 2018). Whether that is possible for martingales and Markov chains is an interesting question for further research.
  • The Wassertein distance used in this paper is the $1$-Wasserstein distance, it would be interesting to obtain convergence rates in the stronger $p$-Wasserstein distance. Work along these lines under different conditions than in this paper can be found in (Bonis 2020), and in the previously mentioned references (Courtade et al. 2019, Zhai 2018). It would be interesting to see if such results can be obtained for martingales and Markov chains.
  • It is interesting to note that, while Theorem 2 is critical to obtain convergence to the central limit theorem for TD learning, the rate of convergence is dominated by the parameter $\delta$ used in the algorithm. It would be interesting to see if the convergence rate result is tight or whether it can be improved through other techniques.
  • Ignoring the $\log$ term in Theorem 4, our result suggests that $2/3$ is a good choice for $\delta$. This is different than the conclusion in (Haque et al. 2023), where their upper bound implies $\delta=0.75$ is the best choice. However, these bounds ignores the constants hidden in the $\mathcal{O}$ which could depends on $\delta$ (see Appendix B). Moreover, what we have is only an upper bound, so a more refined analysis is needed to understand a good choice for $\delta.$
Acknowledgment:
  • I thank Prof. Siva Theja Maguluri, Georgia Tech, for many stimulating discussions and drawing my attention to several references during the course of this work, Prof. Sean Meyn, University of Florida, for pointing out the bias issue with fixed stepsize TD learning and why it cannot be eliminated by averaging, and other discussions, Prof. Mehrdad Moharrami, University of Iowa, for carefully reading an earlier version of the paper and providing many useful comments, and Weichen Wu, CMU, for pointing out a missing factor in Theorem 4 in an earlier version of the paper.
  • Research supported by AFOSR Grant FA9550-24-1-0002, ONR Grant N00014-19-1-2566, and NSF Grants CNS 23-12714, CNS 21-06801, CCF 19-34986, and CCF 22-07547.
Appendix A Properties of $E(||\Delta_{k}||^{2}||$
In both this section and the next, notation that we will use for the constants are just for the particular section of the appendix alone and so we may occasionally reuse some letters that were used for other constants elsewhere in the paper for constants here. We start with the following result from (Srikant and Ying 2019, Theorem 12): there exists a positive integer $k^{*}$ such that for any $k\geq\hat{k},$ we have
$E\left[\|{\Delta}_{k}\|^{2}\right]\leq\kappa_{1}\left(\prod_{j=\hat{k}}^{k-1}a_{j}\right)+\kappa_{2}\sum_{j=\hat{k}}^{k-1}b_{j}\left(\prod_{l=j+1}^{k-1}a_{l}\right),$
where $a_{j}=1-{\kappa_{3}\epsilon_{j}},$ $b_{j}=\kappa_{4}\epsilon_{j}^{2}\log(1/\epsilon_{j}),$ and $\kappa_{1},\kappa_{2},\kappa_{3},\kappa_{4}$ are positive constants. Since $a_{j}\leq\exp(-\kappa_{3}\epsilon_{j}),$ the first term goes to zero exponentially fast. The second term can be upper bounded as follows:
$\kappa_{2}\kappa_{4}\sum_{j=0}^{k-1}\epsilon_{j}^{2}\log\left(\frac{1}{\epsilon_{j}}\right)\exp\left(-\kappa_{3}\sum_{l=j+1}^{k-1}\epsilon_{l}\right)$
$\leq\kappa_{2}\kappa_{4}\delta\log k\sum_{j=0}^{k-1}\epsilon_{j}^{2}\exp\left(-\kappa_{3}\sum_{l=j+1}^{k-1}\epsilon_{l}\right)$
$\leq\kappa_{2}\kappa_{4}\delta\log k\left(\sum_{j=0}^{\lceil k/2\rceil}\exp\left(-\kappa_{3}\sum_{l=j+1}^{k-1}\epsilon_{l}\right)+\sum_{j=\lceil k/2\rceil+1}^{k-1}\epsilon_{j}^{2}\right)$
$=\kappa_{2}\kappa_{4}\delta\log k\left(\sum_{j=0}^{\lceil k/2\rceil}\exp\left(-\kappa_{3}\sum_{l=j+1}^{k-1}(\frac{1}{l+1})^{\delta}\right)+\sum_{j=\lceil k/2\rceil+1}^{k-1}(\frac{1}{j+1})^{2\delta}\right)=\mathcal{O}\left(\frac{\log k}{k^{2\delta-1}}\right).$
Appendix B Properties of $\Upsilon_{j}^{n}$ and $\Phi_{j}^{n}$
Recall the definition of $\Upsilon_{j}^{n}$ and $\Phi_{j}^{n}$ in (12). In this section of the appendix, we examine certain expressions which are shown be either bounded or converging to zero in (Polyak and Juditsky 1992), and derive more explicit bounds for them as a function of $\delta$.
From (Polyak and Juditsky 1992, page 846), we know that $\Phi_{j}^{t}$ can be written as
$\Phi_{j}^{t}=C_{j}^{t}+\bar{A}^{-1}D_{j}^{t},$
where
$||C_{j}^{t}||_{op}\leq\frac{\epsilon_{j}-\epsilon_{j+1}}{\epsilon_{j}}\sum_{i=j}^{t}m_{j}^{i}e^{-\lambda m_{j}^{i}},\qquad D_{j}^{t}=\prod_{l=j}^{t-1}(I-\epsilon_{l}\bar{A}),\mbox{ and }||D_{j}^{t}||_{op}\leq K\exp(-\alpha m_{j}^{t-1}),$
for $t\geq j+1$ and some $K,\alpha,\lambda>0,$ and $m_{j}^{i}=\sum_{k=j}^{i}\epsilon_{k}.$ Since
$\epsilon_{k}=1/(k+1)^{\delta}\geq\mu/(k+1)$
for any $\mu>0$ and any $k\geq\lceil\mu^{1/(1-\delta)}\rceil=:K_{\mu,\delta},$
$m_{j}^{i}=\sum_{k=j}^{i}\epsilon_{k}\geq\mu\sum_{k=j}^{i}\frac{1}{k+1}\geq\mu\log\frac{i+1}{j+1}$
for all $j\geq K_{\mu,\delta}.$ Thus,
$\frac{1}{t}\sum_{j=K_{\mu,\delta}}^{t}||D_{j}^{t}||_{op}\leq K\exp(-\alpha\mu\log\frac{t+1}{j+1})\leq\frac{1}{t}\sum_{j=K_{\mu,\delta}}^{t}K(\frac{j+1}{t+1})^{\alpha\mu}\leq\frac{\kappa}{t}$
for some constant $\kappa.$ Thus,
$\frac{1}{t}\sum_{j=0}^{t}||D_{j}^{t}||_{op}\leq\frac{K\gamma_{\mu,\delta}+\kappa}{t}=\mathcal{O}\left(\frac{1}{t}\right).$
Note that since $\gamma_{\mu,\delta}$ goes to infinity as $\delta\rightarrow 1,$ the rate at which the average of $||D_{j}^{t}||$ goes to zero has a constant that blows up to infinity as $\delta\rightarrow 1.$
Next, let us consider $C_{j}^{i}.$ First we note that, using the fact that $(1+x)^{\delta}\leq 1+\delta x$ for $x\geq 0$ and $\delta\in(0,1),$ we have
$\frac{\epsilon_{j}-\epsilon_{j+1}}{\epsilon_{j}}\leq\frac{\delta}{j+2}.$
Thus,
$||C_{j}^{t}||_{op}\leq\frac{\epsilon_{j}-\epsilon_{j+1}}{\epsilon_{j}}\sum_{i=j}^{t}m_{j}^{i}e^{-\lambda m_{j}^{i}}\leq\frac{1}{j+2}\sum_{i=j}^{t}m_{j}^{i}e^{-\lambda m_{j}^{i}}=\frac{1}{j+2}\sum_{i=j}^{t}m_{j}^{i}e^{-\lambda m_{j}^{i}}\frac{m_{j}^{i}-m_{j}^{i-1}}{\epsilon_{i}}$
$\leq\frac{(j+1)^{\delta}}{j+2}\sum_{i=j}^{t}m_{j}^{i}e^{-\lambda m_{j}^{i}}\frac{m_{j}^{i}-m_{j}^{i-1}}{\epsilon_{i}}\frac{(i+1)^{\delta}}{(j+1)^{\delta}}\leq\frac{1}{(j+2)^{1-\delta}}\sum_{i=j}^{t}m_{j}^{i}e^{-(\lambda-1/\mu)m_{j}^{i}}(m_{j}^{i}-m_{j}^{i-1})$
$=\mathcal{O}\left(\frac{1}{j^{1-\delta}}\right).$
Thus,
$\frac{1}{t}\sum_{j=0}^{t-1}||C_{j}^{t}||_{op}=\mathcal{O}\left(\frac{1}{t^{1-\delta}}\right).$
Finally, since $||\Phi_{j}^{n}||_{op}^{2}\leq 2(||C_{j}^{t}||_{op}^{2}+||D_{j}^{t}||_{op}^{2}),$ and $D_{j}^{t}$ and $C_{j}^{t}$ are uniformly bounded for each fixed $\delta,$ it follows that
$\frac{1}{t}\sum_{j=0}^{t-1}||(\Phi_{j}^{n})^{t}||_{op}^{2}=\mathcal{O}\left(\frac{1}{t^{1-\delta}}\right),$
with a constant that goes to $\infty$ as $\delta\rightarrow 1.$ Thus,
$\frac{1}{t}\sum_{j=0}^{t-1}||(\Upsilon_{j}^{n})^{t}||_{op}^{2}=\mathcal{O}\left(\frac{1}{t^{1-\delta}}\right),$